OpenStudy (anonymous):

Solve. 2x^2-7x+3=2x-1 I just don't understand the steps I need to take.

OpenStudy (anonymous):

x = 1/2 or x = 4

OpenStudy (anonymous):

First step is to get all of your x's on one side

OpenStudy (anonymous):

Which is 2x^-9x+4=0 Right?

OpenStudy (anonymous):